As a hypothetical example, imagine that you were studying a population of Eastern Collared-Lizards (Crotaphytus collaris). The lizards laid 2000 eggs, of which 1000 were eaten by raccoons and skunks. Eight hundred of the surviving juveniles were eaten by foxes and cats. One hundred and eighty of the young adults died during their first hibernation period. What is the indispensable mortality for each stage? If you want to increase the numbers of reproducing adults, which stage should you focus on?
Answer #1

Answer :)

Stages

No of surviving lx

No. of deaths dx

1

2000

1000

2

1000

800

3

200

100

4

100

Indispensable Mortality = [M.S.R. of specific stage] x [Number of adults emerged]

M.S.R. of specific stage = [Mortality in specific stage] / [lx of following stage]

Here lx = Percentage of Number of organisms surviving at the beginning of each interval,

In stage 1, Indispensable Mortality = [M.S.R. of specific stage] x [Number of adults emerged]

M.S.R. of particular stage = 1000/1000

M.S.R. of particular stage = 1

Indispensable Mortality = 1 x 1000

Indispensable Mortality = 1000

In stage 2,

M.S.R. of particular stage = 800 / 200

M.S.R. of particular stage = 4

Indispensable Mortality = 4 x 200

Indispensable Mortality = 800

In stage 3,

M.S.R. of particular stage = 100 / 100

M.S.R. of particular stage = 1

Indispensable Mortality = 1 x 100

Indispensable Mortality = 100

"""""'''""""Please rate the answer""""""""""""